### v2

# Search for spatial coincidence between IceCube neutrinos and gamma-ray bright red dwarfs

Abstract
We search for a spatial coincidence between high energy neutrinos detected by the IceCube neutrino detector and ten red dwarfs which have been observed in gamma-rays. For our analysis, we use the unbinned maximum likelihood method to look for a statistically significant excess. We do not find any such spatial association between any of the red dwarfs and IceCube-detected neutrinos. Therefore, we conclude that none of the gamma-ray bright red dwarfs contribute to the diffuse neutrino flux measured by IceCube.
I Introduction
The origin of majority of the IceCube diffuse neutrino flux in the TeV-PeV energy range Aartsen et al. (2013) is still unknown Halzen (2023). Knowing the sources of the diffuse neutrino background flux will help us unravel a large number of problems in Physics and Astronomy from origin of ultra-high energy cosmic rays to beyond the standard model of particle physics Argüelles et al. (2024). Although, searches by the IceCube collaboration have shown evidence for neutrino emission from some point sources such as NGC 1068, TXS 0506+056, NGC 4151, and PKS 1424+240, the majority of IceCube neutrinos cannot be attributed to any astrophysical sources Abbasi et al. (2021). Searches for spatial coincidence with a plethora of extragalactic sources, which could contribute to the diffuse neutrino flux have been done. A non-exhaustive list of such extra-galactic sources includes AGNs and GRBs Zhou et al. (2021); Hooper et al. (2019); Luo and Zhang (2020); Smith et al. (2021); Li et al. (2022); Abbasi et al. (2011, 2022a); Aartsen et al. (2017); Abbasi et al. (2022b), FRBs Luo and Zhang (2024); Desai (2023); Aartsen et al. (2020a), high energetic events from the Fermi-LAT catalog Li et al. (2022), merging galaxies Bouri et al. (2024), and galaxy catalogs using the 2MASS survey Aartsen et al. (2020b) and WISE-2MASS sample Zhou et al. (2024).
[…]
In a similar vein, a galactic contribution to the all-sky IceCube diffuse neutrino flux cannot be excluded and has been estimated to be about 10-20% Palladino and Vissani (2016); Troitsky (2021). Therefore, searches for coincidences with multiple galactic sources such as supernova remnants, X-ray binaries, pulsars, pulsar wind nebulae, LHAASO and HAWC sources, open clusters have also been carried out Vance et al. (2021); Denton et al. (2017); Abbasi et al. (2022c); Aartsen et al. (2020c); Kovalev et al. (2022); Huang and Li (2022); Li et al. (2022); Kumar et al. (2024); Pasumarti and Desai (2022, 2024). A large number of galactic sources have also been targeted for neutrino astronomy by the previous generation of underground neutrino detectors Ambrosio et al. (2001); Abe et al. (2006); Desai et al. (2008); Desai (2022)
In this work, we look for high energy neutrinos from red dwarfs. Red dwarfs (sometimes also referred to as M-dwarfs) are one of the smallest stars on the main sequence with masses between 0.075-0.5 $M_{\odot}$, having surface temperatures between 2,500-5,000 K Edgeworth (1946), and frequently emit flares with energies between $10^{32}$ and $10^{35}$ ergs with a power-law index between 1.7-2.4 Yang et al. (2017). Energetic outbursts have been detected from red dwarfs at high energies from hard X-rays to TeV gamma-rays Osten et al. (2016); Song and Paglione (2020); Sinitsyna et al. (2019). Most recently, the SHALON TeV gamma-ray telescope detected very high energy gamma-ray emission between 800 GeV and 20 TeV from eight red dwarfs, namely, V388 Cas, V547 Cas, V780 Tau, V962 Tau, V1589 Cyg, GJ 1078, GJ 3684 and GL 851.1 Sinitsyna et al. (2019). Subsequently, these red dwarfs have been proposed as sources of ultra high energy cosmic rays Sinitsyna et al. (2022, 2021). However, no corresponding gamma-ray emission was seen at GeV Huang et al. (2024) or MeV energies Shrivastava et al. (2024)Song and Paglione (2020) reported a $5\sigma$ detection of pulsed GeV $\gamma$-ray emission from another M-dwarf TVLM 513-46546 with Fermi-LAT. A very hard X-ray outburst was detected by the SWIFT-BAT telescope from another M-dwarf, viz DG CVn from 0.3-100 keV in April 2014 Osten et al. (2016), although no coincident gamma-ray emission was seen from Fermi-LAT Loh et al. (2017). Although there are no estimates for high energy neutrino flux emission from red dwarfs, these red dwarfs could generate cosmic rays with energies up to $10^{14}$ eV in their flaring states Sinitsyna et al. (2021). TeV gamma-rays detected from these red dwarfs indicate that these red dwarfs produce flares with total energies of $10^{32}-10^{35}$ ergs Sinitsyna et al. (2022, 2021). If one considers the statistics of M-dwarfs in our galaxy and assuming the frequency of flares to be around 36/year for about 10% of the observed red dwarfs, the total energy released from these flares matches the total energy of cosmic rays in the galactic disk of $\sim 10^{54}$ ergs Sinitysna et al. (2019). During these flares the accelerated protons with energies up to $10^{14}$ eV undergo spallation with the atmosphere outside the star Sinitsyna et al. (2021). Note that similar mechanisms for production of TeV gammma-rays have also been proposed for the Sun Zhou et al. (2017). As argued in  Sinitsyna et al. (2021), these interactions produce pions. Consequently, one would also expect neutrino emission from the aforementioned hadronic interactions along with gamma-rays produced from pion decay. Therefore, these considerations provide enough motivation to search for neutrino emissions to complement recent searches for MeV Shrivastava et al. (2024) and GeV Huang et al. (2024) gamma-rays from these red dwarfs.
Therefore, in this work we look for neutrino emission from the aforementioned 10 red dwarfs using the publicly available muon track data from IceCube. We have previously used this data to look for neutrino emission from pulsars and FRBs Desai (2023); Pasumarti and Desai (2022, 2024). The outline of this manuscript is as follows. The neutrino dataset used for the analysis is discussed in Section II. The analysis and results are discussed in Section III. We conclude in Section IV.
II Dataset
The neutrinos used in this work are taken from the IceCube 10 year muon track data Abbasi et al. (2021). IceCube is a detector located at the South Pole, which detects neutrinos through the Cherenkov radiation emitted by the charged leptons traveling through the ice.
This dataset consists of 1,134,431 track-like neutrino candidates, which were collected between April 2008 (IC-40) and July 2018 (IC 86-V11) from four different phases of the experiment, each having different livetime. Note that for our analysis we have used the augmented dataset analyzed in  Zhou and Beacom (2022) [^1] This dataset is available at https://github.com/beizhouphys/IceCube_data_2008--2018_double_counting_corrected, which corrects for 38 duplicate events found in the IceCube public muon track data due to angular reconstruction error Zhou and Beacom (2022). For each neutrino, the dataset consists of right ascension (RA), declination ($\delta$), error in the position, and reconstructed muon energy.
III Analysis and Results
For our analysis, we are using the unbinned maximum likelihood ratio method (See  Pasumarti and Desai (2022) and references therein, and first proposed in  Braun et al. (2008)). We use neutrinos within a declination of ±$\pm 5^{\circ}$ of the red dwarf. For a dataset containing $N$ events, if $n_{s}$ is the number of signals attributed to the red dwarf, then the likelihood of the entire dataset is given by:
$\mathcal{L}(n_{s})=\prod_{i=1}^{N}\left[\frac{n_{s}}{N}S_{i}+(1-\frac{n_{s}}{N})B_{i}\right]$ (1)
where $S_{i}$ is the signal PDF and $B_{i}$ is the background PDF. Here the signal PDF is given by:
$S_{i}=\frac{1}{2\pi\sigma_{i}^{2}}e^{-(|\theta_{i}-\theta_{s}|)^{2}/2\sigma_{i}^{2}}$ (2)
where $|\theta_{i}-\theta_{s}|$ is the angular distance between the neutrino and the red dwarf, $\sigma_{i}$ is the angular uncertainty in the neutrino position, in radians.The background PDF can be obtained from the solid angle within the declination ($\delta$) of ±$\pm 5^{\circ}$ around each red dwarf and is given as follows Braun et al. (2008):
$B_{i}=\frac{1}{\Omega_{\delta\pm 5^{\circ}}}$ (3)
We then evaluate $\mathcal{L}(n_{s})$ for different values of $n_{s}$, and designate the value of $n_{s}$, for which this likelihood is maximized as $\hat{n}_{s}$. Now to ascertain the significance of the signal, we define the test statistics ($TS$) based on the likelihood ratio, with respect to the null hypothesis in the absence of a signal, ($\mathcal{L}(0)$) as follows:
$TS=2\log\frac{\mathcal{L}(\hat{n}_{s})}{\mathcal{L}(0)}.$ (4)
For the null hypothesis, $TS$ behaves like a $\chi^{2}$ distribution for one degree of freedom Wilks (1938). The detection significance can be obtained from $\sqrt{TS}$. This statistics is also used in gamma-ray astronomy (eg. Manna and Desai (2024)). For a statistically significant detection of $>5\sigma$, TS $>25$. For each of our red dwarfs, we determine $\hat{n}_{s}$ and evaluate $TS$ according to Eq 4. These values of TS, along with the values of $\hat{n}_{s}$ can be found in Table 1. As we can see, none of the red dwarfs show a TS value $>25$. The largest TS value is seen for GJ 1078 with TS of 2.38, whose significance is less than $2\sigma$. Therefore, the observed signal events are consistent with background, and there is no evidence for any spatial association between red dwarfs and IceCube neutrinos, which are observed as muon tracks. Consequently, we calculate the 95% c.l. upper flux limits by calculating the value of $n_{s}$ for which $\Delta(TS)=-3.84$ Zyla et al. (2020), where $\Delta TS=TS-TS_{max}$.
IV Conclusions
In this work, we searched for a spatial coincidence between IceCube neutrinos and 10 gamma-ray bright red dwarfs. The neutrinos used in this work are taken from the publicly released 10 year muon track data observed between 2008-2018, and consisting of 113443 neutrinos. Eight of the 10 red dwarfs were detected at TeV gamma rays by the SHALON telescope. Two others were detected in gamma-rays and X-rays from Fermi-LAT and SWIFT. We analyzed these data using the unbinned maximum likelihood method. A tabular summary of our results for all the red dwarfs can be found in Table 1. All the red dwarfs show a test statistic value of less than two, with significance less than $2\sigma$. Therefore, we conclude that none of the above red dwarfs contribute to the Ice Cube diffuse neutrino flux.
Red dwarf RA () $\delta(^{\circ}$) $\hat{n}_{s}$ $TS$ Upper limit
V388 Cas 15.8326 +62.365 0.0 0.0 16.2
V547 Cas 8.1229 +67.235 0.0 0.0 16.7
V780 Tau 85.107 +24.802 6.00 0.35 29.8
V962 Tau 86.466 +22.879 0.0 0.0 13.8
V1589 Cyg 310.705 +41.38 0.0 0.0 12.7
GJ 1078 80.95 +22.54 20.39 2.38 51.2
GJ 3684 176.77 +70.03 0.0 0.0 14.9
GL 851.1 333.0267 +31.56 0.0 0.0 11.1
TVLM 513-46546 225.284 22.834 0.0 0.0 20.8
DG CVn 202.944 29.28 2.42 0.05 27.2
Table 1: Results from spatial coincidence analysis between IceCube neutrinos and red dwarfs. None of the red dwarfs show a detection significance $>5\sigma$, corresponding to TS$>25$. Therefore, we do not find any spatial association between the red dwarfs and IceCube neutrinos. The last column shows the 95% c.l. upper limit on the number of signal neutrino events.
Acknowledgments
We are grateful to Vibhavasu Pasumarti for making the analysis codes used in  Pasumarti and Desai (2022) publicly available and the anonymous referee for constructive feedback on the manuscript.